A company had 2 million common shares outstanding throughout the last year. Total dividends of $1 million were paid to common stockholders and dividends of $400,000 were paid to preferred stockholders. Net income was $6 million and the tax rate was 40%. The company also had 100,000 options on common stock outstanding throughout the period; the exercise price is $20.00. The average share price over the year was $27.00 and the end year price was $35.00. The diluted earnings per share was closest to:

Group of answer choices

a. $2.74.

b. $2.76.

c. $2.96.
Answer #1

Diluted number of common shares:

Outstanding = 2,000,000

Options exceeding shares on fair price = 100,000 - 100,000*20/27 = 25,925.93 shares

Number of shares = 2,025,925.93

Diluted EPS = (Net Income - Preferred Dividend)/Diluted number of shares

= (6,000,000-400,000)/(2,025,925.93)

= 2.7642 per share

i.e. $2.76 per share

i.e. b is the answer